# Streaming Federated Learning with Markovian Data

### v1

@@ Appendix B Markov chain Preliminaries
The lemma follows by taking $\epsilon=1/4$ as in the definition of mixing time (Definition 3.3). ∎
Appendix C Some useful lemmas
Lemma C.1.
Let $\mu$ and $\nu$ the joint probability distribution of $\left(x_{t}^{(m,k)}\right)_{m\in[M],k\in[K]}$ with initial distribution $\mathcal{Q}$ of $\left(x_{0}^{(m,0)}\right)_{m\in[M]}$ and $\pi=\bigotimes_{m=1}^{M}\pi_{m}$ respectively. Assume that $\mu$ and $\nu$ has well-defined densities $g$ and $h$ with respect to a measure $\rho$ on $\left(\Omega,\mathcal{B}\right)$ as defined in Section 3.2. If Assumptions 3.13.4 and 4.3 hold, then
@@ Appendix C Some useful lemmas
$\mathbb{E}_{\mathcal{Q}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq\frac{2\tau\sigma^{2}}{MK}+\sigma^{2}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.$ (67)
where $P_{\mathcal{Q}}$ and $\mathbb{E}_{\mathcal{Q}}$ denote the data distribution and expectation, respectively, with initial state distribution $\mathcal{Q}$.
@@
We have
$\begin{split}\mathbb{E}_{\mathcal{Q}}&\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla F_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ =&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\int_{\Omega}\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\left(g\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)-h\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\right)\mathrm{d}\rho\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\\ \overset{(i)}{\leq}&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ &+\sigma^{2}\int_{\Omega}\left\lvert g\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)-h\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\right\rvert\mathrm{d}\rho\left(\left(x_{t}^{m}\right)_{m\in[M]}\right)\\ \overset{(ii)}{\leq}&\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]+\sigma^{2}\left\lVert\mu-\nu\right\rVert_{TV}.\end{split}$ (68)
where in $(i)$ we use the uniform bound on $\left\lVert\nabla F_{m}\left(w;x\right)-\nabla F_{m}\left(w\right)\right\rVert^{2}$ (Assumption 4.3) and in $(ii)$, we use the definition of the total variation distance (Definition B.1).
To bound the first term, we utilize [Paulin, 2015, Theorem 3.2], which provides a concentration inequality for the case of stationary Markov chains.This yields
@@
$\begin{split}\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq&\frac{1}{\gamma_{ps}K}\mathbb{E}_{\pi}\left[\left\lVert\frac{1}{M}\sum_{m}\nabla f_{m}\left(w_{t};x\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\\ =&\frac{1}{M^{2}K\gamma_{ps}}\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right].\\ \end{split}$ (69)
Here $\gamma_{ps}$ is the pseudo spectral gap of the transition kernel $P$ of the system-level Markov chain (see [Paulin, 2015, Section 3.1]). With Assumption 3.4[Paulin, 2015, Proposition 3.4] gives $\frac{1}{\gamma_{ps}}\leq 2\tau$. Hence we have:
@@
$\begin{split}\mathbb{E}_{\mathcal{\pi}}\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F\left(w_{t}\right)\right\rVert^{2}\right]\leq&\frac{2\tau}{M^{2}K}\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right]\\ \leq&\frac{2\tau\sigma^{2}}{MK}.\end{split}$ (70)
@@
The last step follows from the fact $\mathbb{E}_{\pi_{m}}\left[\nabla f_{m}\left(w;x\right)\right]=\nabla F_{m}\left(w\right)$. As such, the term $\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right\rVert^{2}\right]$ can be seen as the variance of the sum of $M$ independent variables. Using Assumption 4.3, it then follows that $\mathbb{E}_{\pi}\left[\left\lVert\sum_{m}\left(\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right)\right\rVert^{2}\right]=\sum_{m}\mathbb{E}_{\pi_{m}}\left[\nabla f_{m}\left(w_{t};x\right)-\nabla F_{m}\left(w_{t}\right)\right]\leq M\sigma^{2}$.
Using (51), we also have
$\begin{split}\left\lVert\mu-\nu\right\rVert_{TV}&=\left\lVert P_{\mathcal{Q}}\left(\left(x_{t}^{(m,0)}\right)_{m\in[M]},.\right)-\pi\right\rVert_{TV}\\ &=\left\lVert QP^{Kt}-\pi\right\rVert_{TV}.\end{split}$ (71)
Hence we prove the lemma. ∎
Lemma C.2.
Given Assumption 3.4, with $K\geq\tau$, we have that:
$\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\leq\frac{2\tau}{K\ln 4}+1.$ (72)
Proof.
Since we assume that the product chain is uniformly ergodic, we have that:
$\begin{split}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\leq\left(\frac{1}{4}\right)^{\lfloor Kt/\tau\rfloor}=\left(\frac{1}{4}\right)^{\lfloor K/\tau\rfloor t}.\end{split}$ (73)
By using the geometric sum formula, we have:
$\begin{split}\sum_{t=0}^{T-1}\left(\frac{1}{4}\right)^{\lfloor K/\tau\rfloor t}&\leq\frac{1}{1-(1/4)^{\lfloor K/\tau\rfloor}}\\ &=1+\frac{1}{4^{\lfloor K/\tau\rfloor}-1}\\ &=1+\frac{1}{\exp\left(\lfloor K/\tau\rfloor\ln 4\right)-1}\\ &\leq\frac{1}{\lfloor K/\tau\rfloor\ln 4}+1.\end{split}$ (74)
Here in the last step, we use the inequality $e^{x}\geq 1+x$ for all $x$. Since $K\geq\tau$, we have that $\lfloor K/\tau\rfloor\geq K/(2\tau)$, which proves the lemma. ∎
Appendix D Minibatch SGD

### v2

@@ Appendix B Markov Chain Preliminaries
The lemma follows by taking $\epsilon=1/4$ as in the definition of mixing time (59). ∎
Appendix C Comparison with Existing Work
In Table 1, we compare our results with existing ones in stochastic non-convex optimization with Markovian data in the centralized (or token-passing) setting [3, 15, 14, 17] setting. In the FL setting with i.i.d. data, we include the lower bounds in [42], along with some results in [9]. For completeness, we also report results in the literature of Federated Stochastic Approximation (FSA), which is analogous to stochastic optimization with strongly convex [65] or linear least-square [37] objectives.
Table 1: Comparison of algorithms in terms of communication and sample complexity to reach an $\epsilon$-stationary point for smooth, non-convex objectives.
Additional Constraints: GE: uniform geometric ergodicity (Definition B.10), ST: stationary Markov processes, BG: bounded gradient, BD: bounded domain, FS: finite state space, LT: lower bound on $T$, SS: sample-wise smoothness (Assumption 4.2), BR: bounded Radon–Nikodym derivative (Assumption 3.2), BH: bounded heterogeneity (Assumption 4.4), BHO: bounded heterogeneity around optimum (i.e., the average distance between the optimum of the local objectives and the global objective is bounded, see 37, Page 4), LS: linear least-square objectives, SC: strongly convex global objective..
Algorithm Communication $T$ Sample complexity KTKT
Additional
Constraints
Centralized(1)
@@ Appendix C Comparison with Existing Work
[3, Thm. 3] $\tilde{\mathcal{O}}\!\left(\tau\!\left[\frac{L\Delta_{0}}{\epsilon^{2}}+\frac{L\Delta_{0}\sigma^{2}}{\epsilon^{4}}\right]\right)$ GE
MAG [15, Thm. 4.3] $\tilde{\mathcal{O}}\!\left(\tau\frac{(M+L+G)^{2}G^{2}}{\epsilon^{4}}\right){\textsuperscript{(2)}}$ BG, BD
MC-SGD [17, Thm. 5.3] $\tilde{\mathcal{O}}\!\left(\tau\!\left[\frac{L\Delta_{0}+\sigma^{2}}{\epsilon^{2}}+\frac{(L\Delta_{0}+\sigma^{2})\sigma^{2}}{\epsilon^{4}}\right]\right)$ FS, GE, LT
[14, Thm. 3] $\tilde{\mathcal{O}}\!\left(\tau\frac{L^{2}(1+\left\lVert w^{*}\right\rVert^{2}+\left\lVert w_{0}-w^{*}\right\rVert^{2})}{\epsilon^{4}}\right){\textsuperscript{(3)}}$ FS, GE, LT
FL i.i.d.
Lower bound [42] $\Omega\!\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\Omega\!\left(\frac{L\Delta_{0}\sigma}{M\epsilon^{3}}\right){\textsuperscript{(4)}}$
FedAvg-M [9, Thm. 1] $\mathcal{O}\!\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\mathcal{O}\!\left(\frac{L\Delta_{0}\sigma^{2}}{M\epsilon^{4}}\right)$
FedAvg-M-VR [9, Thm. 2] $\mathcal{O}\!\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\mathcal{O}\!\left(\frac{L\Delta_{0}\sigma}{M\epsilon^{3}}\right){\textsuperscript{(5)}}$ $S$
FSA(6)
FedLSA [37, Cor. 4.5] $\mathcal{O}\!\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right)$ $\tilde{\mathcal{O}}\!\left(\frac{\tau_{\max}}{M\epsilon^{2}}\log\frac{1}{\epsilon}\right)$ GE, ST, BHO, LS
SCAFFLSA [37, Cor. 5.2] $\mathcal{O}\!\left(\log\frac{1}{\epsilon}\right)$ $\mathcal{O}\!\left(\frac{1}{M\epsilon^{2}}\log\frac{1}{\epsilon}\right){\textsuperscript{(7)}}$ BHO, LS
FedHSA [65, Cor. 1] $\tilde{\mathcal{O}}\!\left(\frac{1}{\epsilon}\right)$ $\tilde{\mathcal{O}}\!\left(\frac{\tau_{\max}}{M\epsilon^{2}}\right)$ FS, GE, SC, LT
Ours
Minibatch SGD $\mathcal{O}\!\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\mathcal{O}\!\left(\frac{L\Delta_{0}C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{4}}\right)$ BR
Local SGD $\mathcal{O}\!\left(\frac{L\Delta_{0}}{\epsilon^{2}}\!\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ $\mathcal{O}\!\left(\frac{L\Delta_{0}C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{4}}\!\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ BR, BH, SS
Local SGD-M $\mathcal{O}\!\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right)$ $\mathcal{O}\!\left(\frac{L\Delta_{0}C_{\infty}\sigma^{2}}{\beta\nu_{ps}M\epsilon^{4}}\right)$ BR, SS
  • $\tilde{\mathcal{O}}$ ignores logarithmic constants.
  • (1)
    In the centralized (or token-passing) setting, $\tau$ is the mixing time of the underlying Markov process.
  • (2)
    $M$ and $G$ are upper bounds on the domain diameter and on the gradient norm, respectively.
  • (3)
    $w^{*}$ is a global optimum of the objective. Their proof also requires a projection onto the unknown set that contains all the optima, and hides exponential factors in $\tau$.
  • (4)
    The setting for the lower bound requires other assumptions. We refer to the original paper [42] for a complete list.
  • (5)
    The variance-reduction technique is akin to the STORM algorithm [11], which relies on the sample-wise smoothness assumption (Assumption 4.2) and achieves the sample complexity lower bound in [42]. Extending Local SGD-M with variance reduction to attain this bound is a promising direction.
  • (6)
    This setting focuses on stochastic approximation problems, with strongly monotone or linear operators. The rates presented here denote the complexity to reach an $\epsilon$-optimal solution. We omit problem-specific constants, except $\tau_{\max}$ (the maximum mixing time across all clients) and $M$.
  • (7)
    The authors present only the rate for the i.i.d. case.
We first give a theoretical justification that clarifies why adjusting the step size alone offers limited leverage for controlling the gradient noise term. By a standard decomposition from the $L$-smoothness condition:
[…]
Let $b_{t}=\left\lVert\mathbb{E}_{t}[g_{t}]-\nabla F(w_{t})\right\rVert^{2}$ and $v_{t}=\mathbb{E}_{t}\left[\left\lVert g_{t}-\nabla F(w_{t})\right\rVert^{2}\right]$ be the bias and the variance of the gradient estimate. This bias arises from the Markovian structure of the data, since in the i.i.d. setting, $\mathbb{E}_{t}[g_{t}]=\nabla F(w_{t})$ for all $t$. Observe that the bias is scaled by the step size $\gamma$. Since the term $\left\lVert\nabla F(w_{t})\right\rVert^{2}$ is also scaled by $\gamma$, the bias cannot be directly controlled by the choice of step size without sacrificing the reduction of the objective. Note that by Jensen’s inequality, $b_{t}\leq v_{t}$, hence the bias also contributes to the "variance" term arising in our main results. This leads to a non-vanishing term involving the gradient noise $\sigma^{2}$ in our results as the step size decreases.
We now give additional details of the techniques used in prior works to control this bias, and obtain a vanishing "variance" term. In short, they introduce other hyper-parameters that can be tuned to make the bias vanish, instead of tuning directly the step size directly.
First, [17, 14, 65] decompose the bias term into other terms containing delayed gradient. In detail, in order to bound the bias, instead of taking the conditional expectation up to the current iteration $t$, they condition on a distant past iteration, which is $t$ minus some delay. This delay is later chosen to be scaled with $1/\sqrt{T}$ [17], or with the mixing time $\tau(\epsilon)$ (see Proposition B.11), where $\epsilon$ is then chosen to be scale with the step size [14, 65] (the step size itself is tuned to be at the order of $1/T$). This technique requires a constraint on the minimum number of iterations (or communication rounds) $T$ at the order of $\tilde{\mathcal{O}}(\tau_{\mathrm{mix}})$ (this constraint is referred to as LT in Table 1). As revealed in Table 1, [14] also contains hidden factors that can be exponential in $\tau_{\mathrm{mix}}$. The bound for non-convex functions in [14] (Theorem 3 in the arxiv version) also requires a projection onto the unknown set that contains all the optima.
Next, [3, 15] use minibatch gradient estimates. Precisely, at each iteration, a random batch size is drawn from a truncated geometric distribution. Hence, a new hyper-parameter is introduced, which is the truncation threshold $m$. The authors of [3, 15] show that $b_{t}$ scales inversely with $m$, which is then chosen to be scaled with $\sqrt{T}$.
In the FSA setting, the analysis in [37] employs a blocking technique that sub-samples data sequences to mitigate temporal correlations, using a sub-sampling gap at the order of the maximum mixing time $\tau_{\max}$. Beyond assuming uniform geometric ergodicity, this approach also requires the underlying Markov processes to be stationary. Moreover, their analysis is restricted to linear least-squares problems.
We emphasize that the gradient noise term in our results can be easily made to vanish as $T$ diverges. Consider Minibatch SGD, under the setting of Theorem 4.5, By choosing $K=\mathcal{O}\left(\frac{C_{\infty}T}{\nu_{ps}M}\right)$, then we have:
$\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}$ $\left(\frac{L\Delta_{0}}{T}+\frac{\sigma^{2}}{T}\right),$ (72)
which leads to the sample complexity:
$KT=\mathcal{O}\left(\frac{C_{\infty}(L\Delta_{0}\vee\sigma^{2})^{2}}{\nu_{ps}M\epsilon^{4}}\right).$ (73)
However, the above result shows no improvement in sample complexity compared to those reported in the main paper. We highlight that our complexity results match the best-known rates achieved in the centralized setting [3, 15, 17, 14], scaled by the number of clients $M$ (see Table 1), while avoiding the commonly imposed uniform geometric ergodicity assumption. Instead, we rely on a general mild assumption about the boundedness of the Radon-Nikodym derivative (Assumption 3.2), which allows for a simple change of measures to handle the non-stationarity of Markovian samples when taking conditional expectation. Prior works rely heavily on the exponential mixing property to deal with non-stationarity. The analysis in [3] even yields a dependence on the square of the mixing time [3, Section B.1]. The proof in [15] does not require uniform geometric mixing, but in the other hand relies strongly on the boundedness of the domain. We decide to choose $K$ as independent of both the step size and the number of rounds $T$, because in real-world scenarios, the value of $K$ can be dictated by other factors such as the data-arrival rate and the client’s buffer capacity.
Appendix D Convergence Analysis
We begin this section with the following lemma that bounds the square error of the stochastic gradient estimate computed by $M$ clients, each using $K$ Markovian samples.
Lemma D.1.
Let $q$ be the initial measure of the system-level Markov chain defined in Section 3.2. If Assumptions 3.13.2 and 4.3 hold, then we have:
$\mathbb{E}_{t}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]\leq\frac{4C_{\infty}\sigma^{2}}{\nu_{ps}MK}.$
@@ Appendix D Convergence Analysis
By Markov’s property (27):
$\mathbb{E}_{t}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$
$=\mathbb{E}_{x_{t-1}^{K-1}}\!\left[\left\lVert\frac{1}{MK}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right].$
Hence, by change of measures:
$\mathbb{E}_{t}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$ (74)
$=\mathbb{E}_{\pi}\!\left[\frac{dP\left(x_{t-1}^{K-1},.\right)}{d\pi}\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]$ (75)
@@
$\leq\left\lVert\frac{dP(x_{t-1}^{K-1},.)}{d\pi}\right\rVert_{\pi,\infty}\mathbb{E}_{\pi}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]\quad\textrm{(Holder's inequality)}$ (76)
[…]
We now utilize [43, Theorem 3.7], applied on the stationary system-level Markov process:
@@
$\begin{split}&\mathbb{E}_{\pi}\!\left[\left\lVert\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right\rVert^{2}\right]\\ &=\sum_{j=1}^{d}\mathbb{E}_{\pi}\!\left[\left(\frac{1}{MK}\sum_{m,k}\nabla f_{m}\left(w_{t};x_{t}^{(m,k)}\right)-\nabla F(w_{t})\right)_{j}^{2}\right]\\ &\overset{(i)}{\leq}\frac{4}{K\nu_{ps}}\sum_{j=1}^{d}\mathbb{E}_{\pi}\!\left[\left(\frac{1}{M}\sum_{m}\nabla f_{m}\left(w_{t};x\right)-\nabla F(w_{t})\right)_{j}^{2}\right]\\ &\leq\frac{4}{K\nu_{ps}}\mathbb{E}_{\pi}\!\left[\left\lVert\frac{1}{M}\sum_{m}\nabla f_{m}(w_{t};x)-\nabla F(w_{t})\right\rVert^{2}\right]\\ &\overset{(ii)}{\leq}\frac{4\sigma^{2}}{MK\nu_{ps}}.\end{split}$ (78)
In $(i)$, we apply [43, Theorem 3.7] to each of $d$ functions:
$h_{i}$ $:\Omega\rightarrow\mathbb{R},$
$h_{i}(x_{t}^{k})$ $=\left(\frac{1}{M}\sum_{m}\nabla f_{m}(w_{t};x_{t}^{(m,k)}\right)_{i}.$
@@
In $(ii)$, we use the fact that $\mathbb{E}_{\pi_{m}}\!\left[\nabla f_{m}(w;x)\right]=\nabla F_{m}(w)$ together with the independence between clients’ Markov processes. As such, the term $\mathbb{E}_{\pi}\!\left[\left\lVert 1/M\sum_{m}\nabla f_{m}(w;x)-\nabla F(w)\right\rVert^{2}\right]$ can be seen as the variance of the sum of $M$ independent random variables. Using Assumption 4.3, it follows that:
$\mathbb{E}_{\pi}\!\left[\left\lVert\frac{1}{M}\sum_{m}\nabla f_{m}(w;x)-\nabla F(w)\right\rVert^{2}\right]$ $=\frac{1}{M^{2}}\sum_{m}\mathbb{E}_{\pi_{m}}\!\left[\left\lVert\nabla f_{m}(w;x)-\nabla F_{m}(w)\right\rVert^{2}\right]$
$\leq M\sigma^{2}$
Hence, by plugging (78) back into (77), we prove the lemma.
D.1 Minibatch SGD